0/60km in9.7sec what is the acceleration


1
Expert's answer
2021-09-21T06:33:37-0400
"d=v_0t+\dfrac{1}{2}at^2,""d=0+\dfrac{1}{2}at^2,""a=\dfrac{2d}{t^2}=\dfrac{2\cdot0.60\ km\cdot\dfrac{1000\ m}{1\ km}}{(9.7\ s)^2}=12.75\ \dfrac{m}{s^2}."